solve using any method you choose
-2x+5y=-2
x-3y=3
7x+5y=18
-9y=7x+4

Found 2 solutions by Tobiasz, Alan3354:
Answer by Tobiasz(54)

http://www.algebra.com/tutors/Algebra1_and_2:_How_to_Solve_a_Elimination_Equation.lesson?content_action=show_dev

Answer by Alan3354(69443)

solve using any method you choose
-2x+5y=-2
x-3y=3
I'll assume each pair is a separate problem.
-2x+5y=-2
x-3y=3 --> 2x-6y=6
--------------------
-2x+5y=-2
2x -6y= 6
----------Add
-y = 4
y = -4
-----
x = -9
--------
--------
7x+5y=18
-9y=7x+4 --> -7x-9y=4
Do it like the 1st one.